# https://hertie.ai/

### v1

@@ 3.3 Sparsity constraints enhance class evidence maps
We next compared evidence maps from our model to attribution maps generated with GradCAM [19] on the ViT baseline. As these were multiclass tasks, we only showed class evidence maps from the correctly predicted class. Our class evidence maps, obtained from the convolutional layer before average pooling, clearly highlighted image features relevant to the predicted class (Fig. 2). We noticed that GradCAM on ViT produced cluttered, hard-to-interpret heatmaps. In contrast, the hybrid ResNet-Transformer generated coarser heatmaps due to its large receptive field, while the hybrid BagNet-Transformer provided more localized explanations. The sparse models further refined this by producing sparser heatmaps, focusing decisions on smaller yet relevant retinal regions.
3.4 Evidence maps provide faithful and localized explanations
We quantitatively assessed quality of the explanations by evaluating their precision in identifying DR lesions [14]. Following the International Clinical Diabetic Retinopathy Scale [25], we evaluated three cases: (a) binary evaluation (Fig. 3a), averaging disease-class heatmaps and combining all lesion annotations; (b) severe DR (Fig. 3b), where MAs, HEs, and SEs were combined, and the precision was computed from the severe grade heatmap (c) proliferative DR (Fig. 3c), were all lesions were combined and precision was evaluated from the heatmap from the proliferative grade heatmap. Precision was measured as the proportion of positively activated regions containing lesions [14], using $33\times 3$ 33 non-overlapping patches to match BagNet’s receptive field. For ViT and hybrid FCL models, GradCAM-generated heatmaps were used, and patches were extracted from positively activated regions.
In all cases, the sparse BagNet-Transformer showed considerably higher precision than all other models and outperformed the base BagNet, suggesting that incorporating attention improved both classification and interpretability. The ResNet-Transformer with an explicit class-evidence layer performed worse, likely due to its larger receptive field producing coarser localizations (Fig. 2).
Subsequently, we additionally measured the faithfulness of the explanations by evaluating their ability to identify relevant regions for classification [27]. Using correctly classified test images, we progressively removed top-ranked patches from highlighted heatmap regions and measured the resulting drop in class confidence. For DR detection, the sparse BagNet-Transformer performed best, while standard ViTs performed worst followed by the ResNet-Transformer (Fig. 3e). For AMD severity classification, the sparse ResNet-Transformer outperformed the sparse BagNet-Transformer (Fig. 3f). This may have been due to larger lesion sizes in AMD, favoring CNNs with larger receptive fields. Notably, this aligned with the classification performance, where the ResNet backbone also performed well.
@@ 3.5 Our model enhances interpretability for multi-class tasks
Finally, we visualized class-specific explanations for the dense and sparse BagNet-Transformer. For DR prediction on the Kaggle dataset, both models correctly classified the example (Fig. 4). Heatmaps and class probability distributions were generated in a single forward pass, with the sparse model producing more focused and localized explanations aligned with the predicted class. In other classes, the sparse model showed almost no positive activations, unlike the dense model, which presented a mix of positive and negative evidence. Interestingly, we observed a strong correlation between attention maps and predicted evidence maps, particularly in the sparse model. This suggests that the model effectively captures long-range dependencies in an interpretable way.
@@ 4 Discussion and Conclusion
Here, we introduced the first inherently interpretable hybrid CNN-transformer architecture for medical image classification and applied it to AMD severity classification and DR detection from retinal fundus images. Our model was evaluated on two CNN backbones – ResNet, which already captured global spatial relationships, and BagNet, which relied on aggregation of small local features. The latter was particularly interesting as the SA mechanism could help to address the limited receptive field size of the BagNet. Our transformer module employed dual-resolution convolutional self-attention to capture both global and fine-grained features while preserving strong local inductive biases. Unlike standard models with FCL classifiers, our model used an explicit class evidence layer, producing spatial feature maps that serve as class-evidence heatmaps, enabling direct explanation without post-hoc methods.
Interestingly, and in contrast to sparse BagNet models [14], the interpretability vs. accuracy trade-off was relatively small – all evaluated models performed fairly close to each other with high balanced accuracy and $\kappa$, while the sparse BagNet-Transformer clearly produced the best explanations for DR detection performance, while sparse ResNet-Transformer produced the best explanations for AMD severity classification.
Preliminary experiments revealed that multi-head self-attention increased training time without improving classification, and multi-scale resolution had a limited impact. Following [14], we observed that higher sparsity often led to failures in detecting late stages, especially in DR detection (Fig. 4h), likely due to their underrepresentation in the training set. However, our hybrid CNN-Transformer architecture mitigated this issue, demonstrating its effectiveness in low-data settings. Overall, our work underscores hybrid CNN-Transformer models as a strong alternative to post-hoc ViT explanations, particularly for medical imaging.

### v2

@@ 3.4 Evidence maps provide faithful and localized explanations
We quantitatively assessed the alignment of the explanations with clinical lesion-wise ground truth annotations by evaluating their precision in identifying DR lesions. Following the International Clinical Diabetic Retinopathy Scale [28], we evaluated three cases: (a) binary evaluation (Fig. 3a), averaging disease-class heatmaps and combining all lesion annotations; (b) severe DR (Fig. 3b), where MAs, HEs, and SEs were combined, and the precision was computed from the severe grade heatmap (c) proliferative DR (Fig. 3c), where all lesions were combined and precision was evaluated from the heatmap from the proliferative grade heatmap. Precision was measured as the proportion of positively activated regions containing lesions [17], using $33\times 33$ non-overlapping patches to match BagNet’s receptive field. For ViT and hybrid FCL models, GradCAM-generated heatmaps were used. Patches were extracted from positively activated regions. In all cases, the sparse BagNet-Transformer showed considerably higher precision than all other models and outperformed the base BagNet, suggesting that incorporating attention improved both classification and interpretability. The ResNet-Transformer with an explicit class-evidence layer performed worse, likely due to its larger receptive field producing coarser localizations (Fig. 2).
Subsequently, we additionally measured the faithfulness of the explanations by evaluating their ability to identify relevant regions for classification [30]. Using correctly classified test images, we progressively removed top-ranked patches highlighted in the heatmap and measured the resulting drop in class confidence. For DR detection, the sparse BagNet-Transformer performed best, while standard ViTs performed worst, followed by the ResNet-Transformer (Fig. 3e). In contrast, for AMD severity classification, the hybrid sparse ResNet outperformed the sparse BagNet-Transformer (Fig. 3f), likely due to the larger lesion sizes in AMD, which favor CNNs with larger receptive fields. Notably, this trend was consistent with classification results, where the ResNet backbone also excelled.
@@ 3.5 Our model enhances interpretability for multi-class tasks
Finally, we visualized class-specific explanations for the dense and sparse BagNet-Transformer. For DR prediction on the Kaggle dataset, both models correctly classified the example (Fig. 4). For our hybrid model, heatmaps and class probability distributions were generated in a single forward pass, with the sparse model producing more focused and localized explanations aligned with both the predicted class and clinical ground truth. In contrast, post-hoc explanations required multiple forward passes, increasing the overall inference cost. In other classes, the sparse model showed almost no positive activations, unlike the dense model, which presented a mix of positive and negative evidence. Interestingly, we observed a strong correlation between attention maps and predicted evidence maps, particularly in the sparse model. This suggests that the model effectively captures long-range dependencies in an interpretable way.
@@ 4 Discussion and Conclusion
We introduced the first inherently interpretable hybrid CNN-transformer architecture for medical image classification [^1] Code at https://github.com/kdjoumessi/Self-Explainable-CNN-Transformer, applied to DR detection and AMD severity classification from retinal fundus images. The approach is backbone-agnostic, allowing backbone selection to be guided by disease-specific prior. We evaluated the model with two CNN backbones: ResNet50, which captured global spatial relationships relevant to AMD, and BagNet, which aggregates small local features important for DR detection. The latter is particularly noteworthy, as the SA mechanism helps overcome BagNet’s limited receptive field. Conversely, ResNet’s larger receptive field is better suited for AMD, which involves larger lesions. In both cases, SA enhances the model’s focus on the most relevant features. Our transformer module employs dual-resolution convolutional SA to capture both global and fine-grained features while preserving strong local inductive biases. Unlike standard models with FCL classifiers, our model includes an explicit class evidence layer that produces spatial class-evidence heatmaps, enabling direct interpretability without post-hoc methods.
Interestingly, the interpretability–accuracy trade-off was relatively small, challenging the myth of the accuracy-interpretability tradeoff in self-explainable models [23]. All evaluated models achieved comparable performance, with high balanced accuracy and $\kappa$. Notably, the sparse BagNet-Transformer produced the most informative explanations for DR detection, while the sparse ResNet-Transformer yielded the best explanations for AMD severity classification.
Preliminary experiments showed that multi-head SA increased training time without improving classification, while multi-scale resolution had limited impacts and further increased both the training and inference time—particularly with the BagNet backbone (Tab. 1), due to its larger feature maps and the resulting higher SA computation cost. Following [17], we also observed that higher sparsity often led to missed detection of late-stage DR, likely due to their underrepresentation in the training set (Fig. 4h). However, our hybrid architecture mitigated this issue more effectively, demonstrating robustness in low-data settings. Overall, our findings underscore hybrid CNN-Transformer models as a strong alternative to post-hoc ViT explanations, particularly for medical imaging.